a family is planning to see their home. if they want to be left with $110,000 after paying 12% of the selling price to a realtor as a commission, for how much must they see the house?
Answer by checkley77(12844)

110,000/(1-.12)=110,000/.88=$125,000 selling price.
Proof:
125,000-.12*125,000=125,000-15,000=110,000